OpenStudy (anonymous):

A 40-g ball travelling to the right at 30 cm/s collides head on with an 80-g ball that is at rest. If the collision is perfectly elastic, find the velocity of each ball after collision A. the first ball is going to the right at 10m/s while the other is going to the left at 20m/s B. the first ball is going to the left at 10m/s while the other is going to the right at 20m/s C. the first ball is going to the left at 20 m/s while the other is going to the right at 10 m/s D. the first ball is going to the right at 10 m/s while the other is going to the left at 10 m/s

OpenStudy (anonymous):

For perfectly elastic collision, the kinetic energy and linear momentum are conserved. Can you try to do the problem by yourself now??

OpenStudy (anonymous):

sorry, im not too good with Physics so pardon my indulgence. I tried it but was not getting the answer...

OpenStudy (anonymous):

No problem. Tell me your Conservation of kinetic energy and conservation of linear momentum equation and I will corret them if they are wrong. :)

OpenStudy (anonymous):

*correct

OpenStudy (anonymous):

Check your equations from this link http://en.wikipedia.org/wiki/Elastic_collision#One-dimensional_Newtonian

OpenStudy (anonymous):

Do you mean the energy before collision is equal to the energy after?

OpenStudy (anonymous):

that is, m1u1 + m2u2 = m1v1 + m2v2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now kinetic energy equation. Also What is u2 in our case

OpenStudy (anonymous):

okay, here it is 0 because the block is at rest.

OpenStudy (anonymous):

K.E = 1/2mv^2

OpenStudy (anonymous):

I meant conservation of kinetic energy equation Yes u2 is 0

OpenStudy (anonymous):

1/2mv^2 = mgh ???

OpenStudy (anonymous):

Sorry. It is not right. Think. Initially only one object is moving with velocity v1 and finally the 2 blocks are moving with velocity v' . Now tell me again. If you are still confused, then tell me. I will guide you.

OpenStudy (anonymous):

1/2m1v^2 = (m1+m2)gh ???

OpenStudy (anonymous):

No mgh is potential energy. It should be \[(1/2) m_1v^2 = (1/2)m_1 v_f^2+ (1/2)m_2v_f^2\] Hope you understand it now. Any doubts on how I got this?

OpenStudy (anonymous):

oh yeah..

OpenStudy (anonymous):

Now substitute the given quantities in both equation and solve them.

OpenStudy (anonymous):

Any problem, post it here

OpenStudy (anonymous):

NOTE: Make sure you convert the given quantities in SI units before substituting them.

OpenStudy (anonymous):

yeah, got it...tanx alot

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

please what the answer to this question

OpenStudy (anonymous):

is it a, b, c,or d which

OpenStudy (anonymous):

whats the answer to this equation please